40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 24 17:42:00 2011, Aidan, Update, Green Locking, 15% of end laser sideband power transmitted through cavity 
    Reply  Sat Feb 26 02:33:16 2011, kiwamu, Update, Green Locking, sidebands on beatnote beat_note.png
       Reply  Sun Feb 27 09:43:59 2011, Aidan, Update, Green Locking, sidebands on beatnote 
       Reply  Sun Feb 27 13:09:56 2011, rana, Update, Green Locking, sidebands on beatnote 
Message ID: 4361     Entry time: Sat Feb 26 02:33:16 2011     In reply to: 4351     Reply to this: 4362   4363
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: sidebands on beatnote 

The power ratio of the beatnote signal vs. the 216kHz sideband has been measured.

The measured ratio was -55 dB, which is smaller by about 20 dB than Aidan's estimation.

To confirm this fact we should check the modulation depth of the end PDH somehow.

 

The below is a picture showing the sidebands around the beatnote locked at 66.45 MHz.

Other than the +/-216 kHz sidebands, we can see some funny peaks at +/- 50 kHz and +/-150 kHz

I wonder if they come from the servo oscillation of the MC servo whose UGF is at 24 kHz.  We can check it by unlocking the MC.

beat_note.png

Quote: #4351 by Aidan

So, on the vertex PD, the power of the 80MHz +/-200kHz sidebands should be around sqrt(0.15)*0.05 = 0.02 = 2% of the 80MHz beatnote.

Once we get the green and IR locked to the arm again, we're going to look for the sidebands around the beatnote.

ELOG V3.1.3-